pooja195 (pooja195):

Just out of curiosity can the midpoint formula have fraction coordinates? If so can you please show me an example ? :)

OpenStudy (anonymous):

pick anything

OpenStudy (anonymous):

the average of an even and an odd number is not a whole number

OpenStudy (anonymous):

so for example the midpoint of \((1,3)\) and \((8,10)\) would have fractions

OpenStudy (anonymous):

and of course the midpoint of \((\frac{5}{7},-\frac{8}{3})\) and \((\frac{2}{3},-\frac{5}{4})\) would be nice and ugly

pooja195 (pooja195):

LOL so is it possible to have something like this then? (1/2,5/6) (3,5/4) ? the work you did i already knew but is it EVER possible to have something like this?

OpenStudy (anonymous):

not sure what you are asking you can always find the midpoint of two points, because you can always take the average of any two numbers

pooja195 (pooja195):

lol that answers what i needed thanks :)